OpenStudy (anonymous):

The tail of a kite is 1.5 ft plus twice the length of the kite. Together, the kite and tail are 15 ft in. long. a. Write an expression for the length of the kite and tail together. b. Write 15 ft 6 in. in terms of feet. c. Write and solve an equation to find the length of the tail.

OpenStudy (texaschic101):

(a) t + l = 15 (b) 15 ft 6 in = 15.5 ft or 15 1/2 ft. (c) t = 1.5 + 2l t + l = 15 now sub 1.5 + 2l in for t 1.5 + 2l + l = 15 1.5 + 3l = 15 3l = 15 - 1.5 3l = 13.5 l = 13.5/3 l = 4.5 now sub 4.5 in for l t + l = 15 t + 4.5 = 15 t = 15 - 4.5 t = 10.5 check... t = 1.5 + 2l 10.5 = 1.5 + 2(4.5) 10.5 = 1.5 + 9 10.5 = 10.5 (correct) tail = 10.5 ft